### v1

@@
[…]
Advances in computational power and AI have increased interest in reinforcement learning approaches to inventory management. This paper provides a theoretical foundation for these approaches and investigates the benefits of restricting to policy structures that are well-established by decades of inventory theory. In particular, we prove generalization guarantees for learning several well-known classes of inventory policies, including base-stock and $(s,S)$ policies, by leveraging the celebrated Vapnik-Chervonenkis (VC) theory. We apply the concepts of the Pseudo-dimension and Fat-shattering dimension from VC theory to determine the generalizability of inventory policies, that is, the difference between an inventory policy’s performance on training data and its expected performance on unseen data. We focus on a classical setting without contexts, but allow for an arbitrary distribution over demand sequences and do not make any assumptions such as independence over time. We corroborate our supervised learning results using numerical simulations.
Managerially, our theory and simulations translate to the following insights. First, there is a principle of ?learning less is more? in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies to minimize overfitting errors. Second, the number of parameters in a policy class may not be the correct measure of overfitting error: in fact, the class of policies defined by $T$ time-varying base-stock levels exhibits a generalization error comparable to that of the two-parameter $(s,S)$ policy class. Finally, our research suggests situations in which it could be beneficial to incorporate the concepts of base-stock and inventory position into black-box learning machines, instead of having these machines directly learn the order quantity actions.
@@ 1 Introduction
An emerging alternative approach to inventory management involves learning policies directly from past data. Indeed, under the assumption that demands are uncensored—a common assumption in inventory literature—it becomes possible to accurately evaluate the counterfactual scenario of how any policy would have performed given any historical demand sequence. This allows for the adoption of supervised learning or a ?hindsight? reinforcement learning approach (Sinclair et al. 2023), which has been demonstrated to be computationally feasible in terms of policy differentiation (Alvo et al. 2023) and has even been deployed in practice (Madeka et al. 2022). The latter papers highlight the power of black-box neural networks, which do not require explicit knowledge about the structure of optimal inventory policies to achieve high performance.
By contrast, our paper demonstrates the potential benefits of adhering to policy structures that have been well-established in 70 years of inventory theory, beginning with Arrow et al. (1951). We derive learning theory results and complement them with simulations. We analyze the fundamental classes of base-stock and $(s,S)$ policies, which are optimal in various settings with independent demand over time (e.g., Snyder and Shen 2019). We approach this as a supervised learning problem over demand sequences, in the spirit of Madeka et al. (2022), Alvo et al. (2023), but focus on a classical setting without contexts. We use the celebrated Vapnik-Chervonenkis (VC) theory to bound the learning-theoretic complexity of these policy classes when trained offline on past demand sequences, without assuming they are independent over time as this is often impossible to verify in practice. We provide statistical learning guarantees for this training procedure, which resembles the ?hindsight? reinforcement learning approaches mentioned earlier.
@@
VC theory is a foundational framework in the field of statistical learning theory (Vapnik and Chervonenkis 1971). At a high level, VC theory helps us understand the complexity of a hypothesis class in terms of its ability to generalize to unseen data, which is critical for avoiding overfitting. The core concept of VC theory centers on the VC dimension, along with its generalizations and variations, such as the Pseudo-dimension (Pollard 1984), and the Fat-shattering dimension, also known as the Pseudoγ-dimension (Kearns and Schapire 1994). These dimensions can be used to provide a distribution-free upper bound on the Rademacher complexity, and overall, quantify the capability of a hypothesis class to learn complex patterns. A hypothesis class with a higher dimension generally possesses a greater potential to learn a more complex pattern from data, but may suffer from reduced generalizability due to overfitting, leading to the need for a larger training dataset to ensure good performance of the learned hypothesis.
@@
In the context of inventory management, VC theory can be a powerful tool for establishing rigorous bounds on the generalization error of inventory policy classes, namely, the difference between an inventory policy’s performance on training data and its expected performance on true data. This is useful because, as we will show, naive metrics such as the number of parameters do not necessarily indicate generalizability. We instead use VC theory to indicate generalizability, and corroborate our theoretical results using simulations. All in all, this reveals how effective each of these policy class restrictions is in terms of reducing overfitting, complementing the large body of inventory theory that already characterizes when each of these policy classes are optimal.
@@ 1.2 Outline of Results
We consider inventory replenishment for a single durable good over a finite, discrete time horizon of length $T$. We assume backlogged demand and allow for a deterministic lead time. We define performance by the average inventory cost over the time horizon, which includes fixed costs for each replenishment, backlogging costs, and holding costs. Demands, cost parameters, and policy parameters are normalized so that given any $T$, the performance of any policy on any demand realization is bounded by a constant. The demand realization indicates the demand for the good during each time period, drawn from an arbitrary distribution over sequences of length $T$ (that could exhibit correlations over time). We assume access to $N$ independent and identically distributed (i.i.d.) samples of sequences drawn from the same distribution, and investigate how the uniform generalization error for a policy class $\Pi$ (difference between expected performance on the true distribution, and average empirical performance over the $N$ samples, taking the supremum over policies in $\Pi$) scales with $N$ and $T$. We consider the three different policy classes for $\Pi$: base-stock policies, $(s,S)$ policies, and $(S^{t})$ policies.
@@
Our theoretical results require a problem-specific analysis of inventory trajectories over time, as needed to bound Pseudo-dimension and Pseudoγ-dimension, which to our knowledge is new to both inventory theory and learning theory. We also emphasize that even if one assumes each period to have its own independent demand distribution, results 2. and 3. above were not previous known. In particular, existing works (e.g., Levi et al. 2007, Cheung and Simchi-Levi 2019) have analyzed the sample complexity of $(S^{t})$ policies under the independence assumption (in which case they are optimal) and shown it to grow with $T$, in a related ?multiplicative? setting. Therefore, we believe the non-dependence on $T$ in our additive setting to be quite surprising. We elaborate on these points in Section 4.1.
@@ 2 Further Related Work
Data-driven inventory management: Most papers on data-driven inventory management focus on the Newsvendor model (e.g., Levi et al. 2015, Lin et al. 2022, Besbes and Mouchtaki 2023, Besbes et al. 2023), with some exceptions. Levi et al. (2007) and Cheung and Simchi-Levi (2019) both consider dynamic models and prove sample complexity bounds on base-stock policies under the assumption that demands are independent across time but not necessarily identically distributed. Huh and Rusmevichientong (2009) use stochastic gradient descent to learn inventory policies for demand that is both independent and identically distributed over time. Ban (2020) considers a dynamic model with demand censoring and proves the consistency of a non-parametric estimator of $(s_{t},S_{t})$. Yuan et al. (2021) provide an upper bound on the regret of learning the optimal $(s,S)$ policy.
It is quite surprising that the study of VC-type complexity, even for the simplest classes such as base-stock and ($s,S$) policies, is missing; this is the gap we aim to fill with this paper. It is worth noting that Han et al. (2023) apply existing bounds on the Pseudo-dimension of neural networks directly to solve a Newsvendor problem. Their model is based on the contextual Newsvendor model that was pioneered by Ban and Rudin (2019). By contrast, our analysis focuses on the dynamics of inventory state transitions over time and how this relates to learning-theoretic dimensions. We note that recently, Guan and Mišić (2022) have asked similar questions for optimal stopping, which also considers dynamic decision-making over time but is otherwise unrelated to inventory.
@@ 3 Model and Preliminaries
For the data-driven problem, the training dataset contains $N$ demand samples $\mathbf{d}_{1},\ldots,\mathbf{d}_{N}$ drawn i.i.d. from $\mathcal{D}$, and each of the samples $\mathbf{d}_{i}\triangleq(d_{i}^{t})_{t=1}^{T+L}$ consists of demand realizations from period $1$ to period $T+L$. Note that this assumption is also made in Ban (2020) and Madeka et al. (2022, §7). Here, the $N$ samples can be interpreted as the demands of $N$ similar products over $(T+L)$ periods, or as $N$ records of one product’s demands over $(T+L)$ consecutive periods (e.g., each sample $i=1,\ldots,N$ represents one year and consists of $T+L$ periods).
@@ 3.1 Estimation Error
where the inequality comes from the optimality of $\hat{\pi}$ and the fact that $\mathbb{E}\left[\hat{R}(\pi^{*})-R(\pi^{*})\right]=0$. Note that from an upper bound of the expected estimation error, one can directly derive the sample complexity, i.e., the number of samples required to guarantee a given estimation error with a given high probability. The supremum term in (3.1) is referred to as the (uniform) generalization error (GE) or the representativeness of the policy class $\Pi$ in the literature, which describes the worst-case scenario of overfitting within the policy class. We now formally define the generalization error of the policy class $\Pi$ w.r.t. the function class ${\cal L}(\Pi)$:
@@
It quantifies the generalization ability of an algorithm (or a policy), namely, the gap between the algorithm’s in-sample performance on the training set and its out-of-sample performance. Note that another widely accepted definition of uniform generalization error is $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$, which is commonly used for deriving high-probability sample complexity bounds. The analysis based on this definition is analogous to that by (3.2), expect for a constant factor.
@@ 3.2 Approximation Error
These three classes of policies have been extensively studied in the inventory literature and are widely implemented in practice. We next review several classical results for finite-horizon backlogged inventory models in the stochastic inventory control literature (e.g., Snyder and Shen 2019), demonstrating the conditions under which these policies achieve optimality (i.e., zero approximation error).
@@
  • In settings where $K>0$ and demands are i.i.d. across time, an $(s,S)$ policy with both stationary order-up-to level $S$ and reorder point $s$ is asymptotically optimal (as $T\to\infty$), falling within the class $\Pi_{(s,S)}$ where $H$ and $\underline{H}$ are selected appropriately based on the cost parameters (to be further discussed in Section 4.3).
@@ 3.3 Rademacher Complexity, VC-Dimension, Pseudo-Dimension, and Pseudoγ-Dimension
The empirical Rademacher complexity measures the richness and complexity of the function class $\mathcal{F}(\Pi)$ given the dataset $(\mathbf{d}_{i})_{i=1}^{N}$. However, computing the empirical Rademacher complexity is NP-hard (see, e.g., the second paragraph on page 29 of Mohri et al. 2018). To address this challenge, combinatorial concepts, namely the VC-dimension (Vapnik and Chervonenkis 1971) and Pseudo-dimension (Pollard 1984), are introduced. Compared to Rademacher complexity, VC-dimension and Pseudo-dimension are often relatively easier to bound or estimate, providing a simpler alternative for bounding the expected generalization error. Therefore, sample complexity upper bounds can be directly inferred using the VC-dimension or Pseudo-dimension, allowing researchers to concentrate on the problem-specific task of bounding the VC-dimension or Pseudo-dimension for specific classes of algorithms of interest.
@@ 4.1 Summary of Our Results and Techniques
Our result holds for arbitrary distributions $\mathcal{D}$ over $[0,1]^{T}$, but it is significant even if one assumes independence over time, i.e., $\mathcal{D}$ is a product distribution $\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{T}$. Indeed, the general theory of Guo et al. (2021) implies a sample complexity that grows linearly in the number of marginal distributions $T$ (although technically their result requires discretization or a strong monotonicity condition, we believe with some work their result can be applied here). The independent setting has also been studied in classical data-driven inventory control papers. For example, Levi et al. (2007) and Cheung and Simchi-Levi (2019) show that the number of samples required to make the estimation error $R(\hat{\pi})-R(\pi^{*})$ at most $\epsilon R(\pi^{*})$ is $O(T^{5}/\epsilon^{2})$ and $O(T^{6}/\epsilon^{2})$, respectively. Our result is not directly comparable because they are providing a multiplicative guarantee, along with other differences in the assumptions. For example, the nature of multiplicative guarantees is that one can allow for unbounded demands; however, one now needs to fix cost parameters $b,h$ before considering asymptotics in $N$ and $T$. Nonetheless, the fact that existing bounds even assuming independence grow with $T$ suggests that our $T$-independent sample complexity is quite surprising.

### v2

@@
Advances in computational power and AI have increased interest in reinforcement learning approaches to inventory management. This paper provides a theoretical foundation for these approaches and investigates the benefits of restricting to policy structures that are well-established by decades of inventory theory. In particular, we prove generalization guarantees for learning several well-known classes of inventory policies, including base-stock and $(s,S)$ policies, by leveraging the celebrated Vapnik-Chervonenkis (VC) theory. We apply the concepts of the Pseudo-dimension and Fat-shattering dimension from VC theory to determine the generalization error of inventory policies, that is, the difference between an inventory policy’s performance on training data and its expected performance on unseen data. We focus on a classical setting without contexts, but allow for an arbitrary distribution over demand sequences and do not make any assumptions such as independence over time. We corroborate our supervised learning results using numerical simulations.
Managerially, our theory and simulations translate to the following insights. First, there is a principle of ?learning less is more? in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies to minimize overfitting errors. Second, the number of parameters in a policy class may not be the correct measure of overfitting error: in fact, the class of policies defined by $T$ time-varying base-stock levels exhibits a generalization error an order of magnitude lower than that of the two-parameter $(s,S)$ policy class. Finally, our research suggests situations in which it could be beneficial to incorporate the concepts of base-stock and inventory position into black-box learning machines, instead of having these machines directly learn the order quantity actions.
@@ 1 Introduction
An emerging alternative approach to inventory management involves learning policies directly from past data. Indeed, under the assumption that demands are uncensored—a common assumption in inventory literature—it becomes possible to accurately evaluate the counterfactual scenario of how any policy would have performed given any historical demand sequence. This allows for the adoption of supervised learning or a ?hindsight? reinforcement learning approach (Sinclair et al. 2023), which has been demonstrated to be computationally feasible in terms of policy differentiation (Alvo et al. 2023) and has even been deployed in practice (Madeka et al. 2022). The latter papers highlight the power of black-box neural networks, which do not require explicit knowledge of the structure of optimal inventory policies to achieve high performance.
By contrast, our paper demonstrates the potential benefits of adhering to policy structures that have been well-established in 70 years of inventory theory, beginning with Arrow et al. (1951). Instead of learning arbitrary ordering actions, we learn inventory policies that are defined by the well-established concepts of base-stock levels and reorder points. Not only does this further improve empirical performance, but it is also founded on theoretical guarantees for learning the best policies within these classes. In particular, we analyze the fundamental classes of base-stock and $(s,S)$ policies, which are optimal in various settings with independent demand over time (e.g., Simchi-Levi et al. 2014, Snyder and Shen 2019). We focus on a non-contextual setting where we are given samples from an unknown underlying distribution of demand sequences $(d^{t})_{t=1}^{T}$ (whose entries can be arbitrarily correlated over time $t$) over a $T$-period horizon. We view this parsimonious, context-free setting as an important starting point for understanding the performance of the supervised learning approach described by Madeka et al. (2022) and Alvo et al. (2023). This approach still captures the high-level goal of incorporating arbitrary demand patterns in a model-free way.
We analyze the estimation error, which is the performance gap between the empirical policy learned from training data and the best-in-class policy when evaluated on the true distribution. The estimation error depends on the number of samples, which is the general focus of analysis in data-driven Newsvendor and inventory problems (see Levi et al. 2007). However, papers in this literature assume that the random demands $d^{t}$ are independent over time $t$, and the estimation error is bounded by analyzing the dynamic program for independent demands. Our analysis completely relaxes the independence assumption, motivated by the aforementioned goal of supervised learning over correlated sequences, and in some cases even improves the guarantee on estimation error. Although seemingly paradoxical, eliminating the independence assumption necessitates a completely different analysis based on the celebrated Vapnik-Chervonenkis (VC) theory, which proves useful even for independent demands.
@@
VC theory is a foundational framework in the field of statistical learning theory (Vapnik and Chervonenkis 1971). At a high level, VC theory helps us understand the complexity of a hypothesis class in terms of its ability to generalize to unseen data, which is critical for avoiding overfitting. The core concept of VC theory centers on the VC dimension, along with its generalizations and variations, such as the Pseudo-dimension (Pollard 1984), and the Fat-shattering dimension, also known as the Pseudoγ-dimension (Kearns and Schapire 1994). These dimensions can be used to provide a distribution-free upper bound on the Rademacher complexity, and overall, quantify the capability of a hypothesis class to learn complex patterns. A hypothesis class with a higher dimension generally possesses a greater potential to learn more complex patterns from data. However, it may suffer from reduced generalizability due to overfitting, necessitating a larger training dataset to ensure good performance of the learned hypothesis.
@@
In the context of inventory management, VC theory can be a powerful tool for establishing rigorous bounds on the generalization error of inventory policy classes, namely, the difference between an inventory policy’s performance on training data and its expected performance on true data. This is useful because, as we will show, naive metrics such as the number of parameters do not necessarily indicate generalizability. Bounds on the generalization error can then be directly translated back into bounds on the estimation error of learning best-in-class policies. We instead use VC theory to indicate generalizability, and corroborate our theoretical results using simulations. All in all, this reveals how effective each of these policy class restrictions is in terms of reducing overfitting, complementing the large body of inventory theory that already characterizes when each of these policy classes is optimal.
@@ 1.2 Outline of Results
We consider inventory replenishment for a single durable good over a finite, discrete time horizon of length $T$. We assume backlogged demand and allow for a deterministic lead time. We define performance by the average inventory cost over the time horizon, which includes fixed costs for each replenishment, backlogging costs, and holding costs. Demands, cost parameters, and policy parameters are normalized so that given any $T$, the performance of any policy on any demand realization is bounded by a constant. The realized demand vector $(d^{t})_{t=1}^{T}$ indicates the demand for the good during each time period, drawn from an arbitrary distribution over sequences of length $T$. We assume access to $N$ independent and identically distributed (i.i.d.) samples of sequences drawn from the same distribution, and investigate how the uniform generalization error for a policy class $\Pi$ (difference between expected performance on the true distribution, and average empirical performance over the $N$ samples, taking the supremum over policies in $\Pi$) scales with $N$ and $T$. We consider the three different policy classes for $\Pi$: base-stock policies, $(s,S)$ policies, and $(S^{t})$ policies.
@@
Our theoretical results require a problem-specific analysis of inventory trajectories over time, as needed to bound Pseudo-dimension and Pseudoγ-dimension, which to our knowledge is new to both inventory theory and learning theory. We also emphasize that even if one assumes each period to have its own independent demand distribution, results 2. and 3. above were not previous known. In particular, existing works (e.g., Levi et al. 2007, Cheung and Simchi-Levi 2019) have analyzed the sample complexity of $(S^{t})$ policies under the independence assumption (in which case they are optimal) and shown it to grow with $T$, in a related ?multiplicative? setting. More recently, Qin et al. (2023) have derived results for the same additive setting as ours, and even under independence their sample complexity for $(S^{t})$ policies grows with $T$. Therefore, we believe the non-dependence on $T$ in our additive setting to be quite surprising. We elaborate on these points in Section 4.1.
@@ 2 Further Related Work
Data-driven inventory management: Most papers on data-driven inventory management focus on the Newsvendor model (e.g., Levi et al. 2015, Lin et al. 2022, Besbes and Mouchtaki 2023, Besbes et al. 2023), with some exceptions. Levi et al. (2007), Cheung and Simchi-Levi (2019), and Qin et al. (2023) all consider dynamic models and prove sample complexity bounds on base-stock policies under the assumption that demands are independent across time but not necessarily identically distributed. Huh and Rusmevichientong (2009) use stochastic gradient descent to learn inventory policies for demand that is both independent and identically distributed over time. Ban (2020) considers a dynamic model with demand censoring and proves the consistency of a non-parametric estimator of $(s_{t},S_{t})$. Yuan et al. (2021) provide an upper bound on the regret of learning the optimal $(s,S)$ policy. Fan et al. (2024) demonstrate that one can beat ?best-possible? bounds derived for general MDPs by leveraging the structural properties of inventory dynamics under $(s,S)$ policies.
It is quite surprising that the study of VC-type complexity, even for the simplest classes such as base-stock and ($s,S$) policies, is missing; this is the gap we aim to fill with this paper. It is worth noting that, recently, Chen et al. (2022) and Han et al. (2023) applied existing bounds on the generalization error of neural networks directly to a Newsvendor problem. Their model is based on the contextual Newsvendor model that was pioneered by Ban and Rudin (2019). By contrast, our analysis focuses on the dynamics of inventory state transitions over time and how this relates to learning-theoretic dimensions. We note that recently, Guan and Mišić (2022) have asked similar questions for optimal stopping, which also considers dynamic decision-making over time but is otherwise unrelated to inventory.
@@ 3 Model and Preliminaries
For the data-driven problem, the training dataset contains $N$ demand samples $\mathbf{d}_{1},\ldots,\mathbf{d}_{N}$ drawn i.i.d. from $\mathcal{D}$, and each of the samples $\mathbf{d}_{i}\triangleq(d_{i}^{t})_{t=1}^{T+L}$ consists of demand realizations from period $1$ to period $T+L$. Note that this assumption is also made in Ban (2020) and Madeka et al. (2022, §7). Here, the $N$ samples can be interpreted as the demands of $N$ similar products over $(T+L)$ periods, or as $N$ records of one product’s demands over $(T+L)$ consecutive periods (e.g., each sample $i=1,\ldots,N$ represents one year and consists of $T+L$ periods). The samples can also correspond to the life cycles of the $N$ past generations of a product (e.g. a magazine or video game), whose sales are correlated within cycle (e.g. one generation was popular and had high demands throughout its life cycle) but each cycle can be viewed as an i.i.d. realization.
@@ 3.1 Estimation Error
$\leq\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[R(\hat{\pi})-\hat{R}(\hat{\pi})\right]$
@@
where the first inequality comes from the optimality of $\hat{\pi}$ and the fact that $\mathbb{E}\left[\hat{R}(\pi^{*})-R(\pi^{*})\right]=0$. The supremum term in (3.1) is referred to as the (uniform) generalization error (GE) or the representativeness of the policy class $\Pi$ in the literature, which describes the worst-case scenario of overfitting within the policy class. We now formally define the generalization error of the policy class $\Pi$ w.r.t. the function class ${\cal L}(\Pi)$:
@@
It quantifies the generalization ability of a policy class $\Pi$, namely, the worst-case gap between in-sample and out-of-sample performances of a policy within $\Pi$.
@@
We remark that the literature often focuses on high-probability bounds for the estimation error instead of expectation bounds, in which case the definition of generalization error can be $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$. Regardless, our expectation upper bounds imply high-probability upper bounds after applying standard tools such as McDiarmid’s inequality (e.g., Theorem 26.5 of Shalev-Shwartz and Ben-David 2014), and our analysis of $\sup\limits_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}$ implies upper bounds on $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$ after scaling by a constant factor.
@@ 3.2 Approximation Error
These three classes of policies have been extensively studied in the inventory literature and are widely implemented in practice. We next review several classical results for finite-horizon backlogged inventory models in the stochastic inventory control literature (e.g., Simchi-Levi et al. 2014, Snyder and Shen 2019), demonstrating the conditions under which these policies achieve optimality (i.e., zero approximation error).
@@
  • In settings where $K>0$ and demands are i.i.d. across time, an $(s,S)$ policy with both stationary order-up-to level $S$ and reorder point $s$ is asymptotically optimal (as $T\to\infty$).
@@ 3.3 Rademacher Complexity, VC-Dimension, Pseudo-Dimension, and Pseudoγ-Dimension
The empirical Rademacher complexity measures the richness and complexity of the function class $\mathcal{F}(\Pi)$ given the dataset $(\mathbf{d}_{i})_{i=1}^{N}$. However, computing the empirical Rademacher complexity is NP-hard (see, e.g., the second paragraph on page 29 of Mohri et al. 2018). To address this challenge, combinatorial concepts, namely the VC-dimension (Vapnik and Chervonenkis 1971) and Pseudo-dimension (Pollard 1984), are introduced. Compared to Rademacher complexity, VC-dimension and Pseudo-dimension are often relatively easier to bound or estimate, providing a simpler alternative for bounding the expected generalization error. Therefore, sample complexity upper bounds can be directly inferred using the VC-dimension or Pseudo-dimension, allowing researchers to concentrate on the problem-specific task of bounding the VC-dimension or Pseudo-dimension for specific policy classes of interest.
@@ 4.1 Summary of Our Results and Techniques
Our result holds for arbitrary distributions $\mathcal{D}$ over $[0,1]^{T}$, but it is significant even if one assumes independence over time, i.e., $\mathcal{D}$ is a product distribution $\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{T}$. Indeed, the general theory of Guo et al. (2021) implies a sample complexity that grows linearly in the number of marginal distributions $T$ (although technically their result requires discretization or a strong monotonicity condition, we believe with some work their result can be applied here). The independent setting has also been studied in classical data-driven inventory control papers. For example, Levi et al. (2007) and Cheung and Simchi-Levi (2019) show that the number of samples required to make the estimation error $R(\hat{\pi})-R(\pi^{*})$ at most $\epsilon R(\pi^{*})$ is $O(T^{5}/\epsilon^{2})$ and $O(T^{6}/\epsilon^{2})$, respectively. Our result is not directly comparable because they are providing a multiplicative guarantee, along with other differences in the assumptions. For example, the nature of multiplicative guarantees is that one can allow for unbounded demands; however, one now needs to fix cost parameters $b,h$ before considering asymptotics in $N$ and $T$. Nonetheless, the fact that existing bounds even assuming independence grow with $T$ suggests that our $T$-independent sample complexity is quite surprising. The surprisingness of our result is further echoed in a recent paper by Qin et al. (2023), who study the same additive setting as ours. Their sample complexity of $(S^{t})$ policies is $O(T/\epsilon^{2})$, also growing with $T$. (Their stated sample complexity is $O(T^{3}/\epsilon^{2})$ because their cost function does not divide by $T$.)
Our $T$-independent result also relates to a broad literature on studying dimension-dependence sample complexity bounds in stochastic programming and learning theory. For example, in convex stochastic programming, Liu and Tong (2024) present a dimension-independent sample complexity bound for sample average approximation under a set of assumptions, including strong convexity. In contrast, the loss function of $(S^{t})$ policies in our setting is not even convex in $(S^{t})$ (e.g., Kunnumkal and Topaloglu 2008). For generic Lipschitz-continuous loss functions, Shalev-Shwartz et al. (2009) establish a sample complexity bound that depends on the dimension of the decision variable (i.e., $T$ in our setting). However, we demonstrate that in our specific inventory context, the $T$-dependence is unnecessary. Considering all these aspects, we find our $T$-independent result to be quite surprising.